### v1

# On independence complexes of graph products

Abstract
We study the independence complex for graph products where at least one of the factors is a path as well as the complex es for their induced subgraphs. We calculate the homotopy type of the independence complex of the graphs $P_{n}\times P_{m}$, $P_{n}\boxtimes P_{2}$, $P_{n}\boxtimes P_{3}$ and $P_{n}\boxtimes P_{4}$. We also focus in the independence complexes of the induced subgraphs of $P_{n}\times P_{3}$, $P_{n}\boxtimes P_{2}$, $P_{n}\boxtimes P_{3}$, $P_{n}\boxtimes P_{4}$ and some lexicographic products.
Keywords: Independence complex, grah products, homotopy type
Mathematics Subject Classification: 05C76, 05E45, 55P15
Acknowledgments. This work was supported by UNAM Posdoctoral Program (POSDOC).
1 Introduction
The independence complex is one of the most studied graph complexes in topological combinatorics, it is define as the set of independent sets of a graph. This complex has many applications in other branches in mathematics, for example it has been used to study the extreme Khovanov (co)homology of link diagrams (see [15, 26, 27]). Another way to think the independence complex is to define it as the clique complex of the complement graph, i.e. the simplices are the complete graphs of the complement graph; in a finite metric space the Vietoris-Rips complexes are clique complexes (see for example [2, 3, 6]) and this complexes are used in topological data analysis and geometric group theory (see for example [25]).
In general is not easy to determine the homotopy type of the independence complex. Even for highly symmetric graphs the topology of this complex can be really hard to understand, for example the chessboard complex– which can be seen as the independence complex of the cartesian product of complete graphs– is not well understood and has torsion for many cases (see [14, 18, 29, 30, 31]). For this reason much of the work is concentrated in calculating the homotopy type for specific families of graphs (see [1, 7, 8, 16, 17, 20]) and for graph products of particular graphs (see for example 4, 9, 21, 22, 23). Here we will focus in study the independence complex for graph products where at least one of the factors is a path as well as the complexes for their induced subgraphs. The products $P_{n}\boxtimes P_{m}$ will be our main objective, but we will do some calculations for categorical and lexicographic products.
2 Preliminaries
A simplicial complex $K$ is a family of subsets of an finite set $V(K)$, the vertices of the complex, such that if $\tau\subseteq\sigma$ and $\sigma\in K$, then $\tau\in K$. We will not distinguish between a simplicial complex and its geometric realization.
We are only interested in simple graphs. For a graph $G$, we denote its vertex set as $V(G)$ and its edge set as $E(G)$, the order of the graph is the non-negative integer number $|V(G)|$. For a vertex $v$, $N_{G}(v)=\{u\in V(G):\;uv\in E(G)\}$ is its open neighborhood and $N_{G}[v]=N_{G}(G)\cup\{v\}$ its closed neighborhood. The degree of a vertex will be denoted $d_{G}(v)=|N_{G}(v)|$.
Given a graph $G$, its complement graph is the graph $G^{c}$ with vertex set $V(G)$ and edge set $E(G^{c})=\{\{u,v\}:\;\;\{u,v\}\notin E(G)\}$.
Taking $\underline{n}=\{1,\dots,n\}$, we define the path on $n$ vertices as the graph $P_{n}$ with vertex set $\underline{n}$ and edge set $E(P_{n})=\{ij:\;|i-j|=1\}$; we define the cycle on $n\geq 3$ vertices as the graph $C_{n}$ with vertex set $\underline{n}$ and edge set $E(C_{n})=\{ij:\;|i-j|=1\}\cup\{1n\}$. $K_{n}$ is the complete graph with vertex set $\underline{n}$ and edge set $\{\{i,j\}:\;i\neq j\}$.
For two vertex disjoint graphs $G$ and $H$ their union is the graph $G+H$ with vertex set $V(G)\cup V(H)$ and edge set $E(G)\cup E(H)$. Given two graphs $G$ and $H$, we will consider four different graphs over the vertex set $V(G)\times V(H)$:
  • The categorical product $G\times H$, with edge set
    $\{\{(u_{1},v_{1}),(u_{2},v_{2})\}:\;\{u_{1},u_{2}\}\in E(G)\mbox{ and }\{v_{1},v_{2}\}\in E(H)\}.$
  • The cartesian product $G\oblong H$ with edge set
    $\{\{(u_{1},v),(u_{2},v)\}:\;\{u_{1},u_{2}\}\in E(G)\}\cup\{\{(u,v_{1}),(u,v_{2})\}:\;\{v_{1},v_{2}\}\in E(H)\}.$
  • The strong product $G\boxtimes H$ with edge set $E(G\times H)\cup E(G\oblong H)$.
  • The lexicographic product $G\circ H$ with edge set
    $\{\{(u,v_{1}),(u,v_{2})\}:\;\{v_{1},v_{2}\}\in E(H)\}\cup\{\{(u_{1},v_{1}),(u_{2},v_{2})\}:\;\{u_{1},u_{2}\}\in E(G)\}.$
For all the graph definitions not stated here we follow [10].
A vertex set $\sigma\subseteq V(G)$ is independent if any two vertices $u,v$ do not form an edge in $G$. The independence complex of $G$ is the simplicial complex
$I(G)=\{\sigma\subseteq V(G):\;\sigma\mbox{ is independent}\}$
Lemma 1.
Let $u,v$ be different vertices of a graph $G$:
  1. 1.
    [13] If $N_{G}(u)\subseteq N_{G}(v)$, then $I(G)\simeq I(G-v)$.
  2. 2.
    [26] If $N_{G}[u]\subseteq N_{G}[v]$, then $I(G)\simeq I(G-v)\vee\Sigma I(G-N_{G}[v])$.
Lemma 2.
[9] Let $G$ be a graph with $v$ a vertex of degree $2$ such that $N_{G}(v)=\{v_{1},v_{2}\}$ is an independent set. We take the graph $H$ obtain from $G$ by first erasing $N_{G}(v_{1})\cap N_{G}(v_{2})$ from $G$, and next adding the edges $u_{1}u_{2}$ for all $u_{1}\in N_{G}(v_{1})-N_{G}(v_{2})$ and $u_{2}\in N_{G}(v_{2})-N_{G}(v_{1})$. Then
$I(G-N_{G}(v_{1}))\cup I(G-N_{G}(v_{2}))=I(H)$
$I(G-N_{G}(v))\cap\left(I(G-N_{G}(v_{1}))\cup I(G-N_{G}(v_{2}))\right)=I(H-v_{1}-v_{2})$
For a vertex $v$ such that $N_{G}(v)$ is an independent set, we define its star cluster as the subcomplex of $I(G)$ given by
$SC(v)=\bigcup_{u\in N(v)}I(G-N_{G}(u)).$
Theorem 3.
[5] Let $G$ be a graph and let $v$ be a non-isolated vertex of $G$ which is contained in no triangle. Then
$I(G)\simeq\Sigma(I(G-N_{G}(v))\cap SC(v))$
Theorem 4.
Let $G$ be a graph with $v$ a vertex of degree $2$ such that $N_{G}(v)=\{v_{1},v_{2}\}$ is an independent set. Take $H$ as in Lemma 2, then
$I(G)\simeq\Sigma I(H-v_{1}-v_{2})$
Proof.
The result follows form Lemma 2 and Theorem 3. ∎
Theorem 5.
[11] Let $G$ be a graph and $uv$ an edge. Take $H$ the graph obtained from $G$ by replacing the edge $uv$ by a path of length $4$. Then $I(H)\simeq\Sigma I(G)$.
Theorem 6.
[20]
$I(P_{n})\simeq\left\{\begin{array}[]{cc}\mathbb{S}^{k-1}&\mbox{if }n=3k\\ \ast&\mbox{if }n=3k+1\\ \mathbb{S}^{k}&\mbox{if }n=3k+2\end{array}\right.,\;\;I(C_{n})\simeq\left\{\begin{array}[]{cc}\displaystyle\bigvee_{2}\mathbb{S}^{k-1}&\mbox{if }n=3k\\ \mathbb{S}^{k-1}&\mbox{if }n=3k+1\mbox{ or }n=3k-1\end{array}\right.$
Before ending this section we will define two graph families. We will say a graph $G$ has property $\mathcal{P}$ if $I(G)$ is contractible, has the homotopy type of a wedge of spheres or has the homotopy type of a disjoint union of wedges of spheres. We will say a graph $G$ has property $\mathcal{SP}$ if $I(G)$ is contractible or has the homotopy type of a wedge of spheres. From this, we define the following two families:
$\mathcal{W}=\{G:\mbox{ any induced subgraph of }G\mbox{ has property }\mathcal{P}\}$
$\mathcal{SW}=\{G:\mbox{ any induced subgraph of }G\mbox{ has property }\mathcal{SP}\}$
For example, by Theorem 6, all the paths and cycles are in $\mathcal{SW}$. By the definition of the independence complex is clear that complete graphs are also in $\mathcal{SW}$. It is known that the independence complex of a forest is eihter contractible or has the homotopy type of a sphere [12, see Corollary 6.1], thus all forest are also in $\mathcal{SW}$. All graphs without induced cycles of length divisible by $3$ are also in $\mathcal{SW}$ [19].
3 Independence complexes of categorical
In this section we give a decomposition formula for the homotopy type of the independence complex of $G\times P_{n}$ in terms of joins of the complex $I(G\times P_{2})$. With this we calculate the homotopy type of $I(P_{n}\times P_{m})$. We begin giving a Lemma that was made as a commentary in [4] after the Proposition 3.1 of that article.
Lemma 7.
Let $G$ and $H$ be graphs. Assume there are vertices $u,v$ in $H$ such that $N_{H}(v)\subseteq N_{H}(u)$. Then
$I(G\times H)\simeq I(G\times(H-u))$
Proof.
For any vertex $w$ of $G$, we have that
$N_{G\times H}((w,v))=N_{G}(w)\times N_{H}(v)\subseteq N_{G\times H}((w,u))=N_{G}(w)\times N_{H}(u)$
By Lemma 1 we can erase all the vertices $(w,u)$ without changing the homotopy type. ∎
As immediate consequence we get the following corollary.
Corollary 8.
For any graph $G$, we have that
$I(G\times C_{4})\simeq I(G\times P_{2})$
Theorem 9.
Let $G$ be a graph, then
$I(G\times P_{n})\simeq\left\{\begin{array}[]{cc}I(G\times P_{2})^{*r}&\mbox{ if }n=3r\\ &\mbox{ if }n=3r+1\\ I(G\times P_{2})^{*r+1}&\mbox{ if }n=3r+2\end{array}\right.$
Proof.
For $n=1,2$, the result is clear. For $n\geq 3$, By Lemma 7 we have that
$I(G\times P_{n})\simeq I(G\times(P_{n}-3))\cong I(G\times P_{2})*I(G\times P_{n-3}).$
The results follows by induction. ∎
Corollary 10.
Let $G$ be a bipartite graph, then
$I(G\times P_{n})\simeq\left\{\begin{array}[]{cc}I(G)^{*2r}&\mbox{ if }n=3r\\ &\mbox{ if }n=3r+1\\ I(G)^{*2r+2}&\mbox{ if }n=3r+2\end{array}\right.$
Proof.
Because $G$ is bipartite, we have that $G\times P_{2}\cong 2G$. ∎
As immediate consequence we get the following corollary.
Corollary 11.
Let $G$ be a bipartite graph in $\mathcal{W}$. Then $G\times P_{2}$, then $I(W)$ is in $\mathcal{SW}$.
Proposition 12.
$I(P_{n}\times P_{m})\simeq\left\{\begin{array}[]{cc}\displaystyle\mathbb{S}^{2kr-1}&\mbox{if }n=3k\mbox{ and }m=3r\\ \displaystyle\ast&\mbox{ if }n=3k+1\mbox{ or }m=3r+1\\ \displaystyle\mathbb{S}^{2(k+1)r-1}&\mbox{if }n=3k+2\mbox{ and }m=3r\\ \displaystyle\mathbb{S}^{2(r+1)k-1}&\mbox{if }n=3k\mbox{ and }m=3r+2\\ \displaystyle\mathbb{S}^{2(k+1)(r+1)-1}&\mbox{if }n=3k+2\mbox{ and }m=3r+2\end{array}\right.$
Proof.
Thus result follows from Theorem 6 and Corollary 10. ∎
In [4, Proposition 3.4.] was proved that
$I(C_{m}\times P_{2})\simeq\left\{\begin{array}[]{cc}\displaystyle\bigvee_{4}\mathbb{S}^{4k-1}&\mbox{if }m=6k\\ \mathbb{S}^{4k}&\mbox{if }m=6k+1\\ \mathbb{S}^{4k+1}&\mbox{if }m=6k+r\;\mbox{with }r\in\{2,4\}\\ \mathbb{S}^{4k+1}\vee\mathbb{S}^{4k+1}&\mbox{if }m=6k+3\\ \mathbb{S}^{4k+2}&\mbox{if }m=6k+5\end{array}\right.$
this together with Theorem 9 give us the following corollary.
Corollary 13.
$I(C_{m}\times P_{n})\simeq\left\{\begin{array}[]{cc}\ast&\mbox{ if }n=3r+1\\ \displaystyle\bigvee_{4^{r}}\mathbb{S}^{4kr-1}&\mbox{if }m=6k\mbox{ and }n=3r\\ \mathbb{S}^{4kr+r-1}&\mbox{if }m=6k+1\mbox{ and }n=3r\\ \mathbb{S}^{4kr+2r-1}&\mbox{if }m=6k+l\;\mbox{with }l\in\{2,4\},\mbox{ and }n=3r\\ \displaystyle\bigvee_{2^{r}}\mathbb{S}^{4kr+2r-1}&\mbox{if }m=6k+3\mbox{ and }n=3r\\ \mathbb{S}^{4kr+3r-1}&\mbox{if }m=6k+5\mbox{ and }n=3r\\ \displaystyle\bigvee_{4^{r+1}}\mathbb{S}^{4k(r+1)-1}&\mbox{if }m=6k\mbox{ and }n=3r+2\\ \mathbb{S}^{4k(r+1)+r}&\mbox{if }m=6k+1\mbox{ and }n=3r+2\\ \mathbb{S}^{4k(r+1)+2r+1}&\mbox{if }m=6k+l\;\mbox{with }l\in\{2,4\},\mbox{ and}n=3r+2\\ \displaystyle\bigvee_{4^{r+1}}\mathbb{S}^{4k(r+1)+2r+1}&\mbox{if }m=6k+3\mbox{and }n=3r+2\\ \mathbb{S}^{4k(r+1)+3r+2}&\mbox{if }m=6k+5\mbox{ and }n=3r+2\end{array}\right.$
From what we have done, if $G$ is a cycle or a bipartite graph such that $G^{c}$ is connected, then $I(G\times P_{2})$ is simply connected. This can be generalized for any graph $G$ such that $G^{c}$ is connected.
Theorem 14.
If $G$ is a a graph such that $I(G)$ is connected, then $I(G\times P_{n})$ is simply connected. Even more, for $r\geq 1$
$\mathrm{conn}(I(G\times P_{3r}))\geq r+1\;\;\mbox{and}\;\;\mathrm{conn}(I(G\times P_{3r+2}))\geq r+2.$
Proof.
For $n=1$ the result is clear. If we prove the result is for $n=2$, then, by Theorem 9, the result is true for any $n\geq 3$. Assume $n=2$ and let $T$ be a spanning tree of $G^{c}$. We define $T_{i}$ as the copy of $T$ in $\Delta^{V(G)\times\{i\}}$ and we take $x$ a vertex in $V(G)$. We define
$\mathcal{T}=T_{1}+T_{2}+\{(x,1),(x,2)\}$
a spanning tree of $\left(G\times P_{2}\right)^{c}$. Now we take the free group $H_{\mathcal{T}}$ with generators $\{uv:\;\{u,v\}\in E(\mathcal{T})\}$ and the relations
  • $uv=vu=1$ for all the edges of $\mathcal{T}$.
  • $(ty)(yz)=tz$ if $\{t,y,z\}$ is an independent set.
Then,
$\pi_{1}\left(I(G\times P_{2}),(x,1)\right)\cong H_{\mathcal{T}}$
(see [28, Theorem 7.34]). We will see that $H_{\mathcal{T}}$ is trivial. Take $e$ an edge of $(G\times P_{2})^{c}$:
  • If $e=\{u,v\}$ is an edge of $\mathcal{T}$, then $uv=vu=1$ by definition of the group.
  • If $e=\{(w_{1},1),(w_{2},1)\}$ and it is not in $T_{1}$, we take $y_{0}y_{1}\cdots y_{r}$ the only path form $w_{1}$ to $w_{2}$ in $T_{1}$. Then $(w_{1},1)(w_{2},1)=(y_{1},1)(w_{2},1)$ because $\{(w_{1},1),(w_{2},1),(y_{1},1)\}$ is an independent set and $(w_{1},1)(y_{1},1)=1$. If $r=2$, then $(w_{1},1)(w_{2},1)=1$. Assume $r\geq 3$, then doing the same along the path we get that $(w_{1},1)(w_{2},1)=(y_{1},1)(w_{2},1)=(y_{2},1)(w_{2},1)=\cdots=(y_{r-2},1)(w_{2},1)=((y_{r-2},1)(y_{r-1},1))((y_{r-1},1)(w_{2},1))=1$.
  • If $e=\{(w,1),(w,2)\}$, we take $y_{0}y_{1}\cdots y_{r}$ the only path form $w$ to $x$ in $T_{1}$. For all $i$, we have that
    $(y_{i},1)(y_{i},2)=((y_{i},1)(y_{i},2))((y_{i},2)(y_{i+1},2))=(y_{i},1)(y_{i+1},2)$
    $=((y_{i},1)(y_{i+1},1))((y_{i+1},1)(y_{i+1},2))=(y_{i+1},1)(y_{i+1},2)$
    Therefore $(w,1)(w,2)=(x,1)(x,2)=1$.
  • If $e=\{(w_{1},1),(w_{2},2)\}$, by the previous cases we have that
    $(w_{1},1)(w_{2},2)=((w_{1},1)(w_{2},1))((w_{2},1)(w_{2},2))=1.$
the rest of possibilities are analogous. Therefore $H_{\mathcal{T}}$ is trivial.
The result follows form what we have done and Theorem 9. ∎
The lower bound given in last Theorem is tight, to see this take $C_{5}\times P_{2}\cong C_{10}$, by Theorem 6 $I(C_{5}\times P_{2})\simeq\mathbb{S}^{2}$.
Theorem 15.
$P_{n}\times P_{m}$ is in $\mathcal{SW}$ for all $n$ and $m\leq 4$.
Proof.
For $m=1$, we have that $P_{n}\times P_{1}\cong K_{n}^{c}$ and the result is clear. For $m=2$, we have that $P_{n}\times P_{2}\cong 2P_{n}$ and any induced subgraph is the disjoint union of paths.
For $P_{n}\times P_{3}$ we know it is true for $n\leq 2$. Assume the result is true for $n$ and take $H$ an induced subgraph of $P_{n+1}\times P_{3}$. If neither of the vertices $(1,1),(1,2),(1,3)$ are in $V(H)$, then $H$ is an induced subgraph of a graph isomorphic to $P_{n}\times P_{3}$ and there is nothing to prove.
If $(1,1)$ is a vertex of $H$, then $d_{H}((1,1))\leq 1$ and $I(H)\simeq*$ or $I(H)\simeq\Sigma I(H-N_{H}[(2,2)])$. If $(1,2)$ is not in $V(H)-N_{H}[(2,2)]$, then $H-N_{H}[(2,2)]$ is an induced subgraph of a graph isomorphic to $P_{n}\times P_{3}$. If $(1,2)$ is in $V(H)-N_{H}[(2,2)]$, then $d_{H}((1,2))=d_{H-N_{H}[(2,2)]}((1,2))\leq 2$ and $I(H-N_{H}[(2,2)])\simeq*$ or $I(H-N_{H}[(2,2)])\simeq\Sigma I(H-N_{H}[(2,2)]-N_{H}[(1,2)])$ or
$I(H-N_{H}[(2,2)])\simeq\Sigma I(H-N_{H}[(2,2)]-N_{H}[(1,2)]-(3,2)).$
Regardless of the case, we got that $I(H)$ is either contractible or has the homotopy type of the double suspension of a complex $I(H^{\prime})$ where $H^{\prime}$ is an induced subgraph of a graph isomorphic to $P_{n}\times P_{3}$. If $(1,3)$ is in $H$, we proceed as before. If $(1,2)$ is a vertex of $H$ and $(1,1),(1,3)$ are not in $V(H)$, we proceed as the last case when we assumed $(1,2)$ was a vertex of $H-N_{H}[(2,2)]$.
Lastly, assume $m=4$. We know that the result is true for $n\leq 3$. Assume the result is true for $n$ and take $H$ and induced subgraph of $P_{n+1}\times P_{3}$. If neither of the vertices $(1,1),(1,2),(1,3),(1,4)$ are in $V(H)$, then $H$ is an induced subgraph of a graph isomorphic to $P_{n}\times P_{3}$ and there is nothing to prove. If the vertex $(1,1)$ is in $H$, it has at most degree $1$, then $I(H)$ is contractible or $I(H)\simeq\Sigma I(N_{H}[(2,2)])$. If $(1,4)$ is in $H$, then $d_{H-N_{H}[(2,2)]}((1,4))=d_{H}((1,4))\leq 1$ and $I(H-N_{H}[(2,2)])$ is contractible or
$I(H-N_{H}[(2,2)])\simeq\Sigma I(H-N_{H}[(2,2)]-N_{H}[(2,3)])$
with $H-N_{H}[(2,2)]-N_{H}[(2,3)]$ an induced subgraph of a graph isomorphic to $P_{n}\times P_{4}$. Assume $(1,4)$ is not in $H$. If $(2,4)$ is a vertex of $H$, then $I(H-N_{H}[(2,2)])$ is contractible. Assume $(2,4)$ is not in $H$. If $(1,2)$ is not in $H$, then $H-N_{H}[(2,2)]$ is an induced subgraph of a graph isomorphic to $P_{n}\times P_{4}$. Assume $(1,2)$ is a vertex of $H$, then $d_{H-N_{H}[(2,2)]}((1,2))=d_{H}((1,2))\leq 2$. If $d_{H}((1,2))=0$, then $I(H-N_{H}[(2,2)])$ is contractible. If $d_{H}((1,2))=1$, then
$I(H-N_{H}[(2,2)])\simeq\Sigma I(H-N_{H}[(2,2)]-N_{H}[(2,1)])\mbox{ or }$
$I(H-N_{H}[(2,2)])\simeq\Sigma I(H-N_{H}[(2,2)]-N_{H}[(2,3)]).$
Assume $d_{H}((1,2))=2$. If $(3,2)$ is a vertex of $H$, then
$I(H-N_{H}[(2,2)])\simeq\Sigma I(H-N_{H}[(2,2)]-N_{H}[(1,2)]-(3,2)).$
Assume $(3,2)$ is a vertex of $H$, then
$I(H-N_{H}[(2,2)])\simeq\Sigma I(H-N_{H}[(2,2)]-N_{H}[(1,2)]).$
In either case $I(H)$ is contractible or has the homotopy type of the double suspension of a complex $I(H^{\prime})$ where $H^{\prime}$ is an induced subgraph of a graph isomorphic to $P_{n}\times P_{4}$. If $(1,4)$ is in $H$, we proceed as before. If $(1,2)$ or $(1,3)$ are vertices of $H$ and $(1,1),(1,4)$ are not in $V(H)$, we proceed as the last case when we assumed $(1,2)$ was a vertex of $H-N_{H}[(2,2)]$. ∎
Question 16.
Can Theorem 15 be extended to any $m$?
Before ending this section we want to remark that while the homotopy type of the independence complex of the categorical product of paths was easy to determine, the homotopy type of the independence complexes of $P_{n}\oblong P_{m}$ is only know for $1\leq m\leq 6$ (see [21, 22]).
4 Independence complexes of lexicographic products
Given two graphs $G$ and $H$, the lexicographic product $G\circ H$ is given by taking a copy of $H$ for each vertex of $G$ and adding al the possible edges between two copies if the corresponding vertices are adjacent in $G$. We now will define an analogous general construction for simplicial complexes. Given a simplicial complex $K$ with vertex set $\underline{n}$ and $\underline{L}=\{L_{1},\dots,L_{n}\}$ a family of simplicial complexes we define its polyhedral join as the simplicial complex
$(\underline{L})^{*K}=\bigcup_{\sigma\in K}\left(\bigast_{i\in\sigma}L_{i}\right).$
Then
$I(G\circ H)=\left(\underline{I(H)}\right)^{*I(G)}$
This approach to study independence complexes was used in [9, 23, 24]. In [9] it was proved that for any for any family $\underline{L}$ and a tree $T$:
  • If $T\ncong K_{1,n}$, then
    $(\underline{L})^{*I(T)}\simeq\bigvee_{\sigma\in I(T)}\left(I\left(T-N_{T}[\sigma]\right)*\bigast_{i\in\sigma}L_{i}\right).$
  • If $T\cong K_{1,n}$ and $1$ is the vertex of degree $n$, then
    $(\underline{L})^{*I(T)}=L_{1}\sqcup\bigast_{i=2}^{n+1}L_{i}.$
In [9] it was also showed that for a graph $G$ such that it is a cycle of length at least $5$, then
$(\underline{L})^{*I(G)}\simeq\bigvee_{\sigma\in I(G)}\left(I\left(G-N_{G}[\sigma]\right)*\bigast_{i\in\sigma}L_{i}\right)$
for any any family $\underline{L}$. For a cycle of length $4$ is clear that
$(\underline{L})^{*I(C_{4})}=(L_{1}*L_{3})\sqcup(L_{2}*L_{4}).$
Theorem 17.
Let $H$ be a graph in $\mathcal{SW}$. Take $\mathcal{G}$ the graph family containing trees and cycles. For any $G\in\mathcal{G}$, $G\circ H$ is in $\mathcal{W}$.
Proof.
Take $W$ a induced subgraph of $G\circ H$. Then, there is $G^{\prime}$ an induced subgraph of $G$ and $\underline{H^{\prime}}=\{H_{i_{1}},\dots,H_{i_{j}}\}$ a family of induced subgraphs of $H$, where $j$ is the order of $G^{\prime}$, such that
$I(W)=\left(\underline{H^{\prime}}\right)^{*I(G^{\prime})}.$
Now, $G^{\prime}$ is a cycle or the disjoint union of paths from where we arrive to the result. ∎
The next proposition will allow us to show a result similar to Theorem 17 form more general graphs by taking the suspensions of the complexes.
Proposition 18.
[9] Take $K$ a simplicial complex with vertex set $\underline{n}$ and take $(\underline{J})$ a family of simplicial complexes, then
$(\underline{J})^{*K}\simeq\bigvee_{\sigma\in K}\Sigma\mathrm{lk}(\sigma)\;\ast\bigast_{i\in\sigma}J_{i}.$
Theorem 19.
Let $G$ and $H$ be graphs of order at least two in $\mathcal{SW}$. Then for any induced subgraph $F$ in $G\circ H$, $\Sigma I(F)$ is contractible or has the homotopy type of a wedge of spheres.
Proof.
We proceed as in the proof of Theorem 17 and take the suspension
$\Sigma I(W)=\Sigma\left(\underline{H^{\prime}}\right)^{*I(G^{\prime})}$
and Proposition 18 tell us that this has the homotopy type desired. ∎
Question 20.
Under the hypothesis of Theorem 19, Its is true that $G\circ H$ is in $\mathcal{W}$?
5 Independence complexes of strong products
In this section we work on the independence complex of the strong product. We start with to results that are easy to obtain from the definition of strong product.
Lemma 21.
Let $G$ and $H$ be graphs. If there are vertices $u,v$ in $V(G)$ such that $N_{G}[u]\subseteq N_{G}[v]$, then $N_{G\boxtimes H}[(u,x)]\subseteq N_{G\boxtimes H}[(v,x)]$ for any vertex $x$ in $V(H)$.
Corollary 22.
Let $G$ and $H$ be graphs. If there are vertices $u,v$ in $V(G)$ and $x,y$ in $V(H)$ such that $N_{G}[u]\subseteq N_{G}[v]$ and $N_{H}[x]\subseteq N_{H}[y]$, then
$N_{G\boxtimes H}[(u,x)]\subseteq N_{G\boxtimes H}[(u,y)]\subseteq N_{G\boxtimes H}[(v,y)]$
$N_{G\boxtimes H}[(u,x)]\subseteq N_{G\boxtimes H}[(v,x)]\subseteq N_{G\boxtimes H}[(v,y)]$
Now we work on the independence complex of $P_{n}\boxtimes P_{m}$, which is known as the $n\times m$ King’s graph– given a chessboard of $n\times m$, the simplices are given by a set of kings that cannot attack each other. If $n=1$ or $m=1$, then the graph is the path. For $n=1$ and $m=2$ the graph is $K_{2}$, with the independence complex equal to $\mathbb{S}^{0}$. For $n=m=2$ the graph is $K_{4}$, therefore its independence complex is the wedge of $3$ copies of $\mathbb{S}^{0}$.
Proposition 23.
For $n\geq 4$
$I(P_{n}\boxtimes P_{2})\simeq\Sigma I(P_{n-2}\boxtimes P_{2})\vee\bigvee_{2}\Sigma I(P_{n-3}\boxtimes P_{2})$
Proof.
Taking $G=P_{n}\boxtimes P_{2}$, we have that $N_{G}[(1,1)]=N_{G}[(1,2)]\subseteq N_{G}[(2,2)]$, by Lemma 1 we have that
$I(G)\simeq I(G-(2,2))\vee\Sigma I(G-N_{G}[(2,2)]).$
Now $G-N_{G}[(2,2)]\cong P_{n-3}\boxtimes P_{2}$. We take $G^{\prime}=G-(2,2)$. In this new graph we have that $N_{G^{\prime}}[(1,1)]=N_{G^{\prime}}[(1,2)]\subseteq N_{G^{\prime}}[(2,1)]$, so we can apply Lemma 1 again, observe that $G^{\prime}-N_{G^{\prime}}[(2,1)]=G-N_{G^{\prime}}[(2,2)]$ and that $G^{\prime}-(2,1)\cong K_{2}\sqcup P_{n-2}\boxtimes P_{2}$. ∎
Proposition 24.
For $n\geq 4$
$I(P_{n}\boxtimes P_{3})\simeq\Sigma I(P_{n-2}\boxtimes P_{3})\vee\Sigma I(P_{n-3}\boxtimes P_{3})\vee\Sigma^{2}I(P_{n-3}\boxtimes P_{3})$
Proof.
In $G=P_{n}\boxtimes P_{3}$, we have that $N_{G}[(1,3)]\subseteq N_{G}[(1,2)]$. Thus, by Lemma 1,
$I(G)\simeq(G-(1,2))\vee\Sigma I(G-N_{G}[(1,2)]),$
where $G-N_{G}[(1,2)]\cong P_{n-2}\boxtimes P_{3}$. Now, in $G^{\prime}=G-(1,2)$ we have that $N_{G}[(1,3)]\subseteq N_{G}[(2,2)]$, thus we can apply Lemma 1 again. Now $G^{\prime}-N_{G^{\prime}}[(2,2)]\cong P_{n-3}\boxtimes P_{3}$, and in $G^{\prime\prime}=G^{\prime}-(2,2)$ we have that $N_{G^{\prime\prime}}((1,3))\subseteq N_{G^{\prime\prime}}((3,3)),N_{G^{\prime\prime}}((3,2))$ and $N_{G^{\prime\prime}}((1,1))\subseteq N_{G^{\prime\prime}}((3,1)),N_{G^{\prime\prime}}((3,2))$, thus, applying Lemma 1
$I(G^{\prime\prime})\simeq\Sigma^{2}I(P_{n-3}\boxtimes P_{3}).$
Figure 1: $Q_{4}$
Now we define auxiliary graphs for the case $m=4$. For 00, we take $Q_{0}=K_{1}$ and for $n\geq 1$ we obtain $Q_{n}$ from $P_{n}\boxtimes P_{4}$ by adding a new vertex $x$ and making it adjacent to the vertices $(1,1)$ and $(1,2)$ (see Figure 1).
Proposition 25.
For $n\geq 4$
$I(P_{n}\boxtimes P_{4})\simeq\bigvee_{3}\Sigma^{2}I(P_{n-3}\boxtimes P_{4})\vee\bigvee_{2}\Sigma^{2}I(Q_{n-3})\vee\bigvee_{2}\Sigma^{3}I(Q_{n-4})$
$I(Q_{n})\simeq\Sigma^{2}I(P_{n-2}\boxtimes P_{4})\vee\bigvee_{3}\Sigma^{2}I(Q_{n-3})\vee\bigvee_{2}\Sigma^{3}I(Q_{n-4})$
Proof.
We start with $Q_{n}$. We have that $N_{Q_{n}}[(1,4)]\subseteq N_{Q_{n}}[(1,3)]$, thus we can apply Lemma 1. Now, in $Q_{n}-N_{Q_{n}}[(1,3)]$ the only neighbor of $x$ is $(1,1)$, so by Lemma 1
$I(Q_{n}-N_{Q_{n}}[(1,3)])\simeq I(Q_{n}-N_{Q_{n}}[(1,3)]-(1,2))\cong\Sigma I(P_{n-2}\boxtimes P_{4}).$
Now, in $G=Q_{n}-(1,3)$ we have that $N_{G}[(1,4)]\subseteq N_{G}[(2,3)]$ and we can apply Lemma 1. In $G^{\prime}=G-N_{G}[(2,3)]$ we have that $N_{G^{\prime}}(x)\subseteq N_{G^{\prime}}(2,1)$, thus $I(G^{\prime})\simeq I(G^{\prime}-(2,1))\cong\Sigma I(Q_{n-3})$. Now, in $G^{\prime\prime}=Q_{n}-(1,3)-(2,3)$ by Lemma 1 we have that
$I(G^{\prime\prime})\simeq I(G^{\prime\prime}-(3,4)-(3,3)-(2,2)-(2,1))\simeq\bigvee_{2}\Sigma^{2}I(H)$
where $H=G^{\prime\prime}-(3,4)-(3,3)-(2,2)-(2,1)-(1,2)-(1,1)-(1,4)-(2,4)-x$. In $H$ we have that $N_{H}[(3,1)]\subseteq N_{H}[(3,2)]$, thus
$I(H)\simeq I(H-(3,2))\vee\Sigma I(H-N_{H}[(3,2)])\cong I(Q_{n-3})\vee\Sigma I(Q_{n-4}).$
For $F=P_{n}\boxtimes P_{4}$, $N_{F}[(1,4)]\subseteq N_{F}[(1,3)]$, therefore
$I(F)\simeq I(F-(1,3))\vee\Sigma I(F-N_{F}[(1,3)]).$
In $F-N_{F}[(1,3)]$ the only neighbor of $(1,1)$ is $(2,1)$, thus
$I(F-N_{F}[(1,3)])\simeq I(F-N_{F}[(1,3)]-(3,2)-(3,1))$
where $F-N_{F}[(1,3)]-(3,2)-(3,1)\cong H+K_{2}$ and $H$ the graph we obtain while working with $Q_{n}$, thus
$I(F-N_{F}[(1,3)])\simeq\Sigma I(Q_{n-3})\vee\Sigma^{2}I(Q_{n-4}).$
Now we take $F^{\prime}=F-(1,3)$. We have that $N_{F^{\prime}}[(1,4)]\subseteq N_{F^{\prime}}[(2,3)]$, thus
$I(F^{\prime})\simeq I(F^{\prime}-(2,3))\vee\Sigma I(F^{\prime}-N_{F^{\prime}}[(2,3)]).$
In $F^{\prime}-N_{F^{\prime}}[(2,3)]$, the only neighbor of $(1,1)$ is $(2,1)$, thus
$I(F^{\prime}-N_{F^{\prime}}[(2,3)])\simeq I(F^{\prime}-N_{F^{\prime}}[(2,3)]-(3,1))\cong\Sigma I(P_{n-3}\boxtimes P_{4}).$
Now, in $F^{\prime}-(2,3)$ the only neighbor of $(1,4)$ is $(2,4)$, so we can erase the vertices $(3,4),(3,3)$ without changing the homotopy type– notice that the obtained graph has $2$ connected components, one which is $K_{2}$. We call this new graph $F^{\prime\prime}$. In $F^{\prime\prime}$, $N_{F^{\prime\prime}}[(1,2)]\subseteq N_{F^{\prime\prime}}[(2,2)]$, thus
$I(F^{\prime\prime})\simeq I(F^{\prime\prime}-(2,2))\vee\Sigma I(F^{\prime\prime}-N_{F^{\prime\prime}}[(2,2)]).$
Observe that $F^{\prime\prime}-N_{F^{\prime\prime}}[(2,2)]\cong K_{2}+P_{n-3}\boxtimes P_{4}$. In $F^{\prime\prime\prime}=F^{\prime\prime}-(2,2)$, we have that $N_{F^{\prime\prime\prime}}[(1,2)]\subseteq N_{F^{\prime\prime\prime}}[(2,1)]$, thus
$I(F^{\prime\prime\prime})\simeq I(F^{\prime\prime\prime}-(2,1))\vee\Sigma I(F^{\prime\prime\prime}-N_{F^{\prime\prime\prime}}[(2,1)])$
where $F^{\prime\prime\prime}-N_{F^{\prime\prime\prime}}[(2,1)]\cong K_{2}+P_{n-3}\boxtimes P_{4}$ and $F^{\prime\prime\prime}-(2,1)\cong 2K_{2}+H$ with $H$ the same graph as before. ∎
Using Lemma 1 is easy to obtain the following proposition.
Proposition 26.
$I(P_{n}\boxtimes P_{2})\simeq\left\{\begin{array}[]{cc}\mathbb{S}^{0}&\mbox{if}n=1\\ \displaystyle\bigvee_{3}\mathbb{S}^{0}&\mbox{if }n=2\\ \displaystyle\mathbb{S}^{1}\vee\bigvee_{2}\mathbb{S}^{0}&\mbox{if }n=3\end{array}\right.,\;\;I(P_{n}\boxtimes P_{3})\simeq\left\{\begin{array}[]{cc}\mathbb{S}^{0}&\mbox{if }n=1\\ \displaystyle\mathbb{S}^{1}\vee\bigvee_{2}\mathbb{S}^{0}&\mbox{if }n=2\\ \displaystyle\bigvee_{2}\mathbb{S}^{1}\vee\mathbb{S}^{0}&\mbox{if }n=3\end{array}\right.$
$I(P_{n}\boxtimes P_{4})\simeq\left\{\begin{array}[]{cc}\ast&\mbox{if }n=1\\ \displaystyle\bigvee_{5}\mathbb{S}^{1}&\mbox{if }n=2\\ \displaystyle\bigvee_{2}\mathbb{S}^{2}\vee\bigvee_{3}\mathbb{S}^{1}&\mbox{if }n=3\end{array}\right.,\;\;I(Q_{n})\simeq\left\{\begin{array}[]{cc}\ast&\mbox{if }n=0\\ \displaystyle\mathbb{S}^{1}&\mbox{if }n=1\\ \displaystyle\bigvee_{4}\mathbb{S}^{1}&\mbox{if }n=2\\ \displaystyle\bigvee_{2}\mathbb{S}^{2}&\mbox{if }n=3\end{array}\right.$
From all the work done we obtain the following corollary.
Corollary 27.
For $2\leq m\leq 4$, $I(P_{n}\boxtimes P_{m})$ has the homotopy type of a wedge of spheres.
Now, for any $n$, we define the following polynomials:
  • For $m=2$, $f_{n}(x)=\sum_{i=0}\tilde{\beta}_{i}(I(P_{n}\boxtimes P_{2}))x^{i}$
  • For $m=3$, $g_{n}(x)=\sum_{i=0}\tilde{\beta}_{i}(I(P_{n}\boxtimes P_{3}))x^{i}$
Then $f_{1}(x)=1,\;f_{2}(x)=3,\;f_{3}(x)=x+2,\;g_{1}(x)=1,\;g_{2}(x)=x+2,\;g_{3}(x)=3x+1$ and for all $n\geq 4$
$f_{n}(x)=xf_{n-2}(x)+2xf_{n-3}(x)$
$g_{n}(x)=xf_{n-2}(x)+(x^{2}+x)f_{n-3}(x)$
From this is easy to get the next proposition.
Proposition 28.
The generating function for the $f_{i}^{\prime}s$ is
$F(t)=\frac{2t(t+1)(t+\frac{1}{2})}{1-xt^{2}-2xt^{3}}$
and for the $g_{i}^{\prime}s$ is
$G(t)=\frac{(3x+1)t^{3}+2t^{2}+t}{1-xt^{2}-(x+x^{2})t^{3}}$
Theorem 29.
$P_{n}\boxtimes P_{m}$ is in $\mathcal{SW}$ for any $n$ and $1\leq m\leq 4$.
Proof.
For $m=1$, any induced subgraph is a disjoint union of paths, therefore its independence complex is the join of independence complexes of paths, by Theorem 6 each one is contractible or has the homotopy of a single sphere.
Assume $m=2$ and that the result is false in this case. Let $n$ be the first positive integer for which there is an induced subgraph $H$ of $P_{n}\boxtimes P_{2}$ such that $I(H)$ is not contractible nor has the homotopy type of a wedge of spheres. Let $H$ be a vertex minimal counterexample. Now, by minimality of $n$ either $(1,1)$ or $(1,2)$ is in $H$ and not both, as by Lemma 1 $H$ would not be a counterexample. Because $I(H)$ is not contractible, $(2,1)$ or $(2,2)$ must be in $H$, in any case Lemma 1 tell us that $H$ cannot be a counterexample.
Next we take $m=3$. As before, we suppose the result is false and we take $n$ the first positive integer for which the results fails. By the work done we know $n\geq 3$. We take $H$ a vertex minimal counterexample for $n$, this means $H$ is an induced subgraph of $P_{n}\boxtimes P_{3}$ such that $I(H)$ is not contractible nor has the homotopy type of a wedge of spheres. By minimality of $n$, $H$ must have a vertex of the form $(1,j)$. If $j=1$, because $I(H)$ is not contractible at lest one of the vertices $(1,2),(2,1),(2,2)$ must be in $H$, then $N_{H}[(1,1)]$ is contained in the close neighborhood of any of this vertices. Using Lemma 1 and the minimality of $H$ we see that $H$ could not be a counterexample. Thus $j\neq 1$ and in the same way we can see that $j\neq 3$. With the same argument we can see that $(2,2)$ is not a vertex of $H$. Now, $(3,2)$neither is a vertex of $H$, if is a vertex, then $N_{H}(1,2)\subseteq N_{H}(3,2)$. With the same type of arguments we can see that $\delta(H)\geq 2$ and therefore $(2,1),(2,3),(3,1),(3,3)$ are vertices of $H$. Now $n\geq 4$, otherwise $H\cong P_{5}$ and thus $H$ is not a counterexample.
Now, there are $5$ possibilities:
  1. 1.
    $(4,2)$ is a vertex of $H$ and $(4,1),(4,3)$ are not. In this case $H-N_{H}[(4,2)]$ the disjoin union of two graphs, one of which is isomorphic to $P_{3}$ and $H-(4,2)$ the disjoin union of two graphs, one of which is isomorphic to $P_{5}$, In this graphs the $P_{3}$ is contain in the $P_{5}$ and the other graphs are the same, thus the inclusion
[…]
    is null-homotopic, because $I(P_{3})\simeq\mathbb{S}^{0}$ and $I(P_{5})\simeq\mathbb{S}^{1}$. Therefore
    $I(H)\simeq\Sigma I(H-N_{H}[(4,2)])\vee I(H-(4,2))$
    and $H$ is not a counterexample.
  2. 2.
    $(4,1),(4,2)$ are vertices of $H$ and $(4,3)$ is not. Is $H-N_{H}[(4,1)]$ the disjoin union of two graphs, one of which is isomorphic to $P_{4}$, thus $I(H-N_{H}[(4,1)])\simeq*$ and $I(H)\simeq I(H-(4,1))$. Then $H$ can not be a counterexample.
  3. 3.
    $(4,2),(4,3)$ are vertices of $H$ and $(4,1)$ is not. This case is equivalent to the last case.
  4. 4.
    $(4,1),(4,2),(4,3)$ are vertices of $H$. By Theorem 5 $I(H)\simeq\Sigma I(T)$, where
    $T=(H-(1,2)-(2,1)-(2,3))+(3,1)(3,3).$
    Now, by Lemma 1, $I(T)\simeq I(T-(4,2))\vee\Sigma I(T-N_{T}[(4,2)])$. Where, by minimality of $H$, $I(T-N_{T}[(4,2)])$ is contractible or has the homotopy type of a wedge of spheres. We call $W=T-(4,2)$.
  5. 5.
    $(4,1),(4,3)$ are vertices of $H$ and $(4,2)$ is not. By Theorem 5 $I(H)\simeq\Sigma I(W)$, where
    $W=(H-(1,2)-(2,1)-(2,3))+(3,1)(3,3).$
Now, for $W$, for either of the last two cases, if $n=4$, then $I(W)\simeq*$ and $H$ is not a counterexample. Assume $n\geq 5$, then we have that
$I(W)\simeq\Sigma\left(I(W-N_{W}[(3,3)]-N_{W}[(4,3)])\cup I(W-N_{W}[(3,3)]-N_{W}[(3,1)])\right).$
Now, it is not hard to see that
$I(W-N_{W}[(3,3)]-N_{W}[(4,3)])\cup I(W-N_{W}[(3,3)]-N_{W}[(3,1)])=I(Q)$
where $Q=(W-(3,3)-(3,1)-(4,3))+(4,1)(5,3)$ if $(5,3)$ is a vertex of $W$ and if not is not a vertex of $W$ then $Q=W-(3,3)-(3,1)-(4,3)$. In either of case, $Q$ is isomorphic to a proper induced subgraph of $H$. Therefore there is no counterexample.
Lastly we take $m=4$. As before, we suppose the result is false and we take $n$ the first positive integer for which the results fails. By the work done we know that $n\geq 4$. Let $H$ be a vertex minimal counterexample. If $(1,1)$ is in $H$, then it must have degree at least $1$, if not $I(H)$ would be contractible. Let $x$ be a neighbor of $(1,1)$, then $N_{H}[(1,1)]\subseteq N_{H}[x]$ and, by Lemma 1,
$I(H)\simeq I(H-x)\vee\Sigma I(H-N_{H}[x])$
which would imply that $H$ is not a counterexample. Therefore $(1,1)$ is not a vertex of $H$. Similarly $(1,4)$ is not a vertex of $H$. Now, at least one of $(1,2)$ and $(1,3)$ is in $H$, otherwise $H$ would be an induced subgraph of a graph isomorphic to $P_{n}\boxtimes P_{4}$. First assume both vertices are in $V(H)$, then $(2,2)$ and $(2,3)$ are not in $V(H)$. Both $(1,2)$ and $(1,3)$ must have degree $2$, otherwise $I(H)\cong\Sigma I(H-N_{H}[(1,2)])$ or $I(H)\cong\Sigma I(H-N_{H}[(1,3)])$ and $H$ would not be a counterexample. Then $H$ is like Figure 2(a), where we know that the solid vertices and edges are in $H$ and the other may be. By Theorem 4 $I(H)\simeq\Sigma I(W)$ where $W$ is the graph in Figure 2(b) obtained from $H-(1,2)-(1,3)-(2,4)$ by adding the edges $\{(2,1),(3,b)\}$ if $(3,b)$ is in $H$. If $(3,1)$ or $(3,4)$ are not in $W$, the $W$ is an induced subgraph of a graph isomorphic to $P_{n}\boxtimes P_{4}$. Then $(3,1)$ and $(3,4)$ are in $W$.
If $(3,2)$ is in $V(W)$, then $I(W)\simeq I(W-(3,2))\vee\Sigma I(W-N_{W}[(3,2)])$. Now, $W-N_{W}[(3,2)]$ is an induced subgraph of a graph isomorphic to $P_{n}\boxtimes P_{4}$, thus we only are interested in $W-(3,2)$. If $(3,3)$ is in $W$, then $I(W-(3,3))\simeq I(W-(3,2)-(3,3))\vee\Sigma I(W-(3,2)-N_{W}[(3,3)])$ and again, we are only interested in $W-(3,2)-(3,3)$. So we can assume that neither $(3,2)$ nor $(3,3)$ are in $W$. If both $(4,1)$ and $(4,2)$ are in $W$, then $I(W)\simeq I(W-(4,2))\vee\Sigma I(W-N_{W}[(4,2)])$, where as before we are only interested in $W-(4,2)$. Thus we can assume that only one of $(4,1)$ and $(4,2)$ is in $W$. By the same argument we can assume that only one of $(4,3)$ and $(4,4)$ is in $W$. Up to isomorphism there are only three possible cases shown in Figures 2(c), 2(e) and 2(e).
(a) $H$
(b) $W$
(c) $W$ case (i)
(d) $W$ case (ii)
(e) $W$ case (iii)
Figure 2: Auxiliary graphs used in Theorem 29 (I).
Case (i): By Theorem 4, we have that $I(W)\simeq\Sigma I(W-N_{W}[(2,1)])$, where $W-N_{W}[(2,1)]$ is an induced subgraph of a graph isomorphic to $P_{n}\boxtimes P_{4}$. Therefore this case is not possible.
Case (ii): By Theorem 5, $I(W)\simeq\Sigma I(R)$ where $R$ is obtained from $W-N_{W}[(2,1)]$ by adding the edge $\{(4,1),(4,4)\}$ (see Figure 3(a)). With arguments similar at what we have done, we can assume that only one vertex of $(5,1)$ and $(5,2)$ is in $R$, the same for $(5,3)$ and $(5,4)$. Therefore there are four possible cases shown in Figures 3(b), 3(c) and 3(d). Case (a): This case is not possible because $R$ would be an induced subgraph of a graph isomorphic to $P_{n-3}\boxtimes P_{4}$. Case (b): In this case we have that $N_{R}((4,1))\subseteq N_{R}((5,3))$ and $N_{R}((4,4))\subseteq N_{R}((5,2))$, thus $I(R)\simeq I(R-(5,2)-(5,3))$. This can not happen as $R-(5,2)-(5,3)$ is the disjoint union of an edge and an induced subgraph of a graph isomorphic to $P_{n-5}\boxtimes P_{4}$. Case (c): With the same arguments as before we can assume that only one vertex of $(6,1)$ and $(6,2)$ is in $R$, the same for $(6,3)$ and $(6,4)$. Thus there are four cases shown in Figures 3(e), 3(f), 3(g) and 3(h). The first two cases are not possible because $R$ would be an induced subgraph of a graph isomorphic to $P_{n-3}\boxtimes P_{4}$. In other two cases, by Theorem 4, $I(R)\simeq\Sigma I(R-(4,1)-(4,4)-(5,1))$, but $R-(4,1)-(4,4)-(5,1)$ is an induced subgraph of a graph isomorphic to $P_{n-4}\boxtimes P_{4}$. From all that we get that case (ii) is not possible.
(a) $R$
(b) $R$ case (a)
(c) $R$ case (b)
(d) $R$ case (c)
(e) Case (c1)
(f) Case (c2)
(g) Case (c3)
(h) Case (c4)
Figure 3: Auxiliary graphs used in Theorem 29 (II).
Case (iii): As we have done before, we can assume that only one vertex of $(5,1)$ and $(5,2)$ is in $R$, the same for $(5,3)$ and $(5,4)$. There are four possible cases shown in Figures 4(a), 4(b), 4(c) and 4(d). In any case, by Theorem 5, we can add the edge $\{(4,1),(4,3)\}$ to obtain a graph $T$ such that $I(W)\simeq\Sigma I(T)$. In the first case we have that $N_{T}((4,1))\subseteq N_{T}((5,3))$, thus $I(H)\simeq I(T-(5,3))$ by Lemma 1. But $T-(5,3)$ is an induced subgraph of a graph isomorphic to $P_{n-3}\boxtimes P_{4}$, therefore this case is not possible. In the next two case we have that $T$ is an induced subgraph of a graph isomorphic to $P_{n-3}\boxtimes P_{4}$, therefores these case are not possible. If $n=5$, then $T\cong P_{4}$ and $I(T)\simeq*$. Then $n\geq 6$. By Theorem 4, $I(T)\simeq\Sigma I(W^{\prime})$, with $W^{\prime}$a graph like the graph in Figure 2(b) and we are again in case (iii). If $n=6$, then $W^{\prime}\cong P_{5}$ (see Figure 2(e)) and $I(W^{\prime})\simeq\mathbb{S}^{1}$. If $n=7$, then $W^{\prime}\cong P_{7}$ and $I(W^{\prime})\simeq*$. If $n=8$ using Theorem 4 we get a graph similar to the graph in Figure 2(b). We can keep doing this, but the graph is finite, thus we see that this case is not possible.
(a) Case (iii(a))
(b) Case (iii(b))
(c) Case (iii(c))
(d) Case (iii(d))
(e) $H$
(f) $F$
Figure 4: Auxiliary graphs used in Theorem 29 (III).
Lastly, we assume that $(1,2)$ is in $V(H)$ and $(1,3)$ is not. If $(2,2)$ is a vertex of $H$, by Lemma 1 we have that
$I(H)\simeq I(H-(2,2))\vee I(H-N_{H}[(2,2)])$
and we are only interested in $H-(2,2)$. So we can assume that $(2,2)$ is not a vertex of $H$. Then $(2,1)$ and $(2,3)$ are in $V(H)$, otherwise $(1,2)$ would have degree $1$ or 00 and $H$ would not be a counterexample. If $(2,4)$ is a vertex of $H$, then
$I(H)\simeq I(H-(2,3))\vee\Sigma I(H-N_{H}[(2,3)])\;\;\mbox{ and }$
$I(H-(2,3))\simeq\Sigma I(H-(2,3)-N_{H}[(2,1)]).$
This would imply that $H$ is not a counterexample. Thus $(2,4)$ is not a vertex of $H$. If $(3,2)$ is in $V(H)$, then $I(H)\simeq I(H-(3,2))$ by Lemma 1 because $N_{H}((1,2))\subseteq N_{H}((3,2))$. Thus we can assume that $(3,2)$ is not in $V(H)$ (see Figure 4(e)). If $(3,1)$ is not in $H$, then $d_{H}((2,1))=1$ and $I(H)\simeq\Sigma I(H-N_{H}[(1,2)])$. This would imply that $H$ is not a counterexample, thus $(3,1)$ is a vertex of $H$. If $d_{H}((1,3))=1$, then $I(H)\simeq\Sigma I(H-N_{H}[(1,2)])$ and $H$ would not be a counterexample. If $(3,3)$ and $(3,4)$ are in $V(H)$, then
$I(H)\simeq I(H-(3,3))\vee\Sigma I(H-N_{H}[(3,3)])$
and $I(H-N_{H}[(3,3)])$ would have the homotopy type of a wedge of spheres or would be contractible. Therefore we can assume that only one of $(3,3)$ and $(3,4)$ is in $V(H)$. If $(3,4)$ is in $V(H)$, then we are in the case(ii) with $H$ like the graph in Figure 2(d). Then $(3,4)$ is not in $V(H)$. Taking $F$ as the graph obtained from $H-N_{H}[(1,2)]$ by adding the edge $\{(3,1),(3,3)\}$ (see Figure 4(f)), we have that $I(H)\simeq\Sigma I(F)$ by Theorem 5. As we have done before, we can assume that only one vertex of $(4,3)$ and $(4,4)$ is in $F$. If $(4,2)$ is a vertex of $F$, then $N_{F}[(3,1)]\subseteq N_{F}[(4,2)]$ and
$I(F)\simeq I(F-(4,2))\vee\Sigma I(F-N_{F}[(4,2)])$
where we are only interested in $I(F-(4,2))$. Thus we can assume that $(4,2)$ is not a vertex of $F$. There are two possible cases. If $(4,4)$ is in $V(F)$, then $F$ would be an induced subgraph of a graph isomorphic to $P_{n-2}\boxtimes P_{4}$. Assume that $(4,3)$ is in $V(F)$. If $n=4$, then $F\cong P_{4}$ and $I(F)\simeq*$. Thus $n\geq 5$. By Theorem 4 $I(F)\simeq\Sigma I(F^{\prime})$ where $F^{\prime}$ is like the graph in Figure 2(b) and we are in a case already done. Therefore there is no counter example for the result. ∎
Question 30.
Can Theorem 29 be extended to any $m$?
Theorem 31.
Let $m$ be a positive integer such that for any $n$, $P_{n}\boxtimes P_{m}$ is in $\mathcal{SW}$, then $I(P_{n}\boxtimes P_{m+k})$ has the homotopy type of a wedge of spheres for $1\leq k\leq 4$.
Proof.
We take $G=P_{n}\boxtimes P_{m+k}$, with $k=1,2$. Then, for all $j$ in $\underline{m+l}$ we have that
$N_{G}[(j,1)]\subseteq N_{G}[(j,2)]\mbox{ and }N_{W}[(j,1)]\subseteq N_{W}[(j,2)]$
for any induced subgraph $W$ such that $(j,1)$ and $(j,2)$ are in $V(W)$. This last fact together with Lemma 1 will allow us to decompose $I(G)$ into a wedge of spaces we know have the homotopy type of a wedge of spheres. Taking the vertex set $S=\{(1,2),(2,2).\dots(n,2)\}$, the idea is to remove in each step a vertex or a vertex and its closed neighborhood. As fist step we have that
$I(G)\simeq I(G-(1,2))\vee\Sigma I(G-N_{G}[(1,2)])$
Next, we have that
$I(G-(1,2))\simeq I(G-(1,2)-(2,2))\vee\Sigma I(G-(1,2)-N_{G}[(2,2)])$
$I(G-N_{G}[(1,2)])\simeq\Sigma I(G-N_{G}[(1,2)]-(3,2))\vee\Sigma I(G-N_{G}[(1,2)]-N_{G}[(3,2)])$
We keep doing this until we get a family of complexes $I(W_{1}),\dots,I(W_{p})$ such that none of them have any vertex of $S$. To each of these complexes we can associate a sequence $s^{j}=(s_{1},s_{2},\dots,s_{r_{j}})$ such that $s_{i}=1,2$ and $s_{1}+\cdots+s_{r_{j}}=n$ or $s_{1}+\cdots+s_{r_{j}}=n+1$, where:
  1. 1.
    For $1$, $s_{1}$ tell us if we get the complex form $I(G-(1,2))$ or $I(G-N_{G}[(1,2)])$.
  2. 2.
    For $i\geq 2$, $s_{i}$ tell us if in the step $i$ we removed $(s_{1}+\cdots+s_{i-1}+1,2)$ or its closed neighborhood.
For each $j=1,\dots,p$, we take $t_{j}$ equal to the number of entries of $s^{j}$ which are equal to $2$. Therefore
$I(G)\simeq\bigvee_{j=1}^{p}\Sigma^{t_{j}}I(W_{j})$
Now, each $W_{j}$ is the disjoint union of a induced subgraph of $P_{n}\boxtimes P_{m}$ and paths– we are taking $P_{0}=\emptyset$. From this we get the result.
Take $G=P_{n}\boxtimes P_{m+k}$, with $k=3,4$ and do the same as before. Then
$I(G)\simeq\bigvee_{j=1}^{p}\Sigma^{t_{j}}I(W_{j})$
Now, we take $R=\{(1,m+k-1),(2,m+k-1).\dots(n,m+k-1)\}$. As before
$N_{G}[(j,m+k)]\subseteq N_{G}[(j,m+k-1)]\mbox{ and }N_{W}[(j,m+k)]\subseteq N_{W}[(j,m+k-1)]$
for any induced subgraph $W$ such that $(j,m+k)$ and $(j,m+k-1)$ are in $V(W)$. For each $W_{j}$ we do the same process as before on the vertex set $R$. ∎
Corollary 32.
For any $n$ and $1\leq m\leq 4$, $I(P_{n}\boxtimes P_{m})$ has the homotopy type of a wedge of spheres.
Proof.
This is a direct consequence of Theorem 29 and Proposition 31. ∎
Question 33.
Is the homotopy type of $I(P_{n}\boxtimes P_{m})$ always a wedge of spheres for all $n,m$?
A positive answer to Question 30 will give us a positive answer to last question by Theorem 29.